# Machine learning surrogates for efficient hydrologic modeling: Insights from stochastic simulations of managed aquifer recharge

### v1

@@ Abstract
Process-based hydrologic models are invaluable tools for understanding the terrestrial water cycle and addressing modern water resources problems. However, many hydrologic models are computationally expensive and, depending on the resolution and scale, simulations can take on the order of hours to days to complete. While techniques such as uncertainty quantification and optimization have become valuable tools for supporting management decisions, these analyses typically require hundreds of model simulations, which are too computationally expensive to perform with a process-based hydrologic model. To address this gap, we propose a hybrid modeling workflow in which a process-based model is used to generate an initial set of simulations and a machine learning (ML) surrogate model is then trained to perform the remaining simulations required for downstream analysis. As a case study, we apply this workflow to simulations of variably saturated groundwater flow at a prospective managed aquifer recharge (MAR) site. We compare the accuracy and computational efficiency of several ML architectures, including deep convolutional networks, recurrent neural networks, vision transformers, and networks with Fourier transforms. Our results demonstrate that ML surrogate models can achieve under 10% mean absolute percentage error and yield order-of-magnitude runtime savings over processed-based models. We also offer practical recommendations for training hydrologic surrogate models, including implementing data normalization to improve accuracy, using a normalized loss function to improve training stability and downsampling input features to decrease memory requirements.
@@ 1 Introduction
Future water security hinges on effective water resources management. Process-based hydrologic and reactive transport models, which simulate water flow and solute transport (Xu and Singh, 2004; Devia et al., 2015; Steefel et al., 2015), are frequently used to forecast future changes in the hydrologic cycle and to optimize water management strategies. When paired with uncertainty quantification and parameter estimation methods, these models can support decision-making in diverse scenarios such as flood forecasting, contaminant transport, nutrient availability and managed aquifer recharge (MAR). However, these techniques often require hundreds of model simulations in which uncertain input parameters are varied randomly from one simulation to the next. While each individual simulation is deterministic – producing one fixed output for a given set of input parameters – the ensemble of simulations is stochastic because of the random variations in model input. When performed at the scale and resolution required for locally relevant policy decisions (Wood et al., 2011), these stochastic approaches result in large computational loads, even on state-of-the-art supercomputers.
@@
While these early studies highlight the potential for ML surrogate models to accelerate hydrologic simulations, many questions remain for practitioners interested in applying ML emulators to water resources problems. For example, how many process-based simulations are required to train a high-fidelity surrogate model? Given the high dimensionality associated with spatial uncertainty, does the minimum training set size change when varying spatially heterogeneous input parameters across stochastic simulations? How does the choice of loss function impact performance? How important is data normalization? And, given the computational expense associated with training many ML architectures, when is it more computationally efficient to perform simulations exclusively with a process-based hydrologic code instead of an ML surrogate?
In this paper, we investigate these questions and present practical considerations for using ML surrogates to perform stochastic hydrologic simulations. As a case study, we focus on 3D simulations of variably saturated groundwater flow during managed aquifer recharge (MAR), the intentional replenishment of depleted aquifers through recharge basins and injection wells. We investigate seven different surrogate model architectures and evaluate the accuracy with which each one can reproduce output from a parallel, integrated hydrologic code (ParFlow-CLM). Importantly, we explore the impacts of spatial uncertainty on surrogate performance by varying parameters both spatially across the domain and between simulations. We then evaluate the memory and runtime required for each surrogate architecture as compared to the process-based hydrologic code. We also identify a range of implementation techniques that improve model performance, including data normalization, loss normalization and downsampling of high-dimensional inputs using an autoencoder.
@@ 2 Material and Methods
In this section, we formalize the hybrid modeling workflow (§2.1) and introduce the case study (§2.2). Next, we describe the process-based hydrologic code (§2.3) and the surrogate model architectures (§2.4). We then describe data preprocessing (§2.5) and detail the experimental setup for training and evaluating each ML surrogate (§2.6). We conclude with a summary of experiments, enumerating the ML models trained to assess our approach (§2.7).
2.1 Proposed hybrid workflow
The goal of this hybrid workflow is to harness the fast output rate of ML models to efficiently generate a large volume of simulations for downstream hydrologic analysis. Our task is to execute $N$ simulations, generating $N$ simulation outputs from $N$ distinct sets of simulation inputs. As illustrated in Figure 1, we employ a traditional processed-based hydrologic model (ParFlow-CLM) to generate a training set $\mathcal{D}_{\text{train}}$ comprising simulation input–output pairs. Subsequently, an ML surrogate model is trained on these examples and used to generate simulation outputs for the remaining $N-|\mathcal{D}_{\text{train}}|$ simulations, where $|\mathcal{D}_{\text{train}}|$ is the number of examples in $\mathcal{D}_{\text{train}}$. If the ML surrogate provides comparable accuracy to the process-based model, the combined group of $N$ simulations can then be used for further analyses, such as uncertainty quantification, parameter estimation and resource optimization.
@@ 2.2 The case study: Managed aquifer recharge site
We apply this hybrid workflow to a single case study in which hydrologic simulations are used to evaluate a recharge efficiency at a prospective managed aquifer recharge (MAR) site. MAR is a water management strategy whereby land managers deliberately flood parcels of land or inject water into wells to replenish underlying aquifers. Recent work has shown that MAR is a valuable tool for mitigating groundwater depletion in many regions of the globe (Dillon et al., 2019; Sprenger et al., 2017), but evaluating prospective recharge sites remains challenging as subsurface heterogeneity can significantly impact recharge rates (Knight et al., 2022; Perzan et al., 2023). Hydrologic and reactive transport simulations are effective tools for quantifying outcomes of interest at prospective MAR sites, such as changes in groundwater storage (e.g., Ganot et al., 2018) or the fate of subsurface contaminants (e.g., Perzan and Maher, 2024). However, generating these simulations comes with a notable computational expense, as the time required for a single simulation can range from hours to days depending on model scale and resolution. Thus, MAR site selection is a logical case study that could benefit from a computationally efficient hybrid modeling workflow.
@@
The domain is parameterized through a stochastic geophysical-geostatistical workflow (Perzan et al., 2023). Because subsurface heterogeneity can exert a dominant control on recharge efficiency, we parameterize the domain using a subsurface electrical resistivity model derived from a towed transient electromagnetic (tTEM) survey of the site. The tTEM system is a hydrogeophysical tool that can map meter-scale variations in electrical resistivity, a property directly related to sediment lithology. We first generate 600 gridded realizations of the tTEM-derived resistivity model using sequential Gaussian simulation, a stochastic technique for populating a rectilinear grid with a random variable. We then compare vertical profiles of electrical resistivity to collocated profiles of sediment type collected through cone penetration tests. Using the bootstrapping procedure of Knight et al. (2018), we transform each resistivity realization to a realization of coarse fraction, a metric that describes the fraction of coarse-grained (sand and gravel) sediment within each grid cell. The remainder of sediment within the cell is assumed to be fine-grained (silt and clay). For example, a cell with a coarse fraction of 0.5 is 50% coarse-grained material (either 100% sand, 100% gravel, or some combination of the two) and 50% fine-grained material (some combination of silt and clay). To assign hydrogeologic parameters (hydraulic conductivity, porosity, etc.) to each coarse fraction realization, we assume that only two types of sediment exist within a single realization — coarse-grained and fine-grained — and that each cell is a mix of the two types. We then sample from broad distributions to assign parameter values to each end member and use a weighted averaging procedure to calculate values for cells that are a mix of the two end members. The input parameter distributions are derived from a mix of literature and site-specific data. For more information about these parameter values, we refer the reader to the more complete description provided by Perzan et al. (2023).
@@ 2.3 The high-fidelity process-based model: ParFlow-CLM
In the first phase of our proposed framework (§2.1), a process-based hydrologic model is used to generate an initial set of simulations, which are then used to train an ML surrogate model. In this case study, we perform these process-based simulations using ParFlow-CLM. ParFlow-CLM couples the integrated hydrologic code ParFlow with the Common Land Model (CLM) to solve partial differential equations that describe water and energy fluxes over and beneath the land surface. ParFlow solves the Richards’ equation for variably saturated subsurface flow using a Newton-Krylov approach and the shallow-wave equations for overland flow using a kinematic wave approximation. CLM, meanwhile, uses a mass transfer approach to compute water and energy fluxes at the land surface (Maxwell and Miller, 2005; Kollet and Maxwell, 2008), including irrigation, evapotranspiration and precipitation. To accurately simulate the impacts of diurnal temperature and radiation fluctuations on water fluxes across the land surface, we run ParFlow-CLM using an hourly time step in Stages 2 and 3. For a detailed description of the governing equations, we refer readers to previous work describing ParFlow-CLM (Jones and Woodward, 2001; Ashby and Falgout, 1996; Kollet and Maxwell, 2006; Maxwell, 2013; Maxwell and Miller, 2005; Kollet and Maxwell, 2008).
@@ 2.4 The machine learning models
In the second phase of this hybrid framework (§2.1), we train an ML surrogate on the initial set of simulations generated by the process-based model. In total, we evaluate the performance of seven different machine learning surrogate model architectures. The architectures include recurrent and highway networks developed for spatiotemporal forecasting (Wang et al., 2018), vision transformers (Dosovitskiy et al., 2021), convolutional nerual networks and U-shaped networks paired with Fourier transforms (Wen et al., 2022; Zhou et al., 2022; Kang et al., 2023). We selected these architectures because some of them have been successfully applied in previous hydrologic surrogate modeling studies (Tran et al., 2021; Maxwell et al., 2021; Wen et al., 2022), while others are some of the most common architectures used for spatiotemporal forecasting in related fields like computer vision.
The ML surrogate architectures can be categorized into three types based on their method of prediction (Table 1, Figure 2). First, the Non-sequential models emulate the entire simulation in a single forward pass (i.e., the emulator produces 3D pressure fields for all output time steps in one run of the model). By contrast, the One-step models emulate one output time step (producing a single 3D pressure field) per forward pass. These models then use the output from the previous time step to perform another forward pass, repeating until all 3D pressure fields have been generated. Finally, the Recurrent models are similar to the one-step models in that they emulates one output time step per forward pass. However, they use memory cells to retain hidden states between forward passes, so that information can be stored between output time steps. As discussed above, the output frequency varies between model stages (see 2.2), which results in a different shape input tensor for each model stage. Because most ML architectures require a consistent shape input tensor across all training examples, we train a separate surrogate for each model stage. The ML models are summarized below, though complete details are provided in the Supplementary material (Section S1.1).
@@
We experiment with three non-sequential models. Since non-sequential models produce 3D pressure fields for all output time steps in a single forward past, they are four-dimensional; i.e., they receive four-dimensional inputs — x, y, z and time — and generate four-dimensional outputs (3D transient pressure fields). The non-sequential models are:
@@
    ViT4d. The ViT4d is a vision transformer (Dosovitskiy et al., 2021) that uses a four-dimensional convolutional layer to transforms the input tensor into a sequence of four-dimensional patches. This sequence of patches is then fed into a self-attention module (Vaswani et al., 2017) that calculates the importance of each patch relative to every other patch, allowing the model to learn both spatial and temporal dependencies regardless of their distance from each other.
@@
One-step models, unlike non-sequential models, receive three-dimensional inputs (x, y and z) and predict one time step at a time. The output of the previous time step is reused as input for the next time step and this rolling forward process continues until the final time step. Note that these models receive no explicit temporal information during training, such as day of year or time step size. These one-step models include:
@@
    PredRNN++. This recurrent model is an improved predictive recurrent neural network (PredRNN++), which was originally developed by Wang et al. (2018). The PredRNN++ architecture (Figure S2) pairs causal long short-term memory networks (causal LSTMs) with a gradient highway units (GHUs). Causal LSTMs contain a spatiotemporal memory mechanism designed to improve short-term spatial correlations, while GHUs connect future outputs with distant inputs for improved long-term modeling capabilities. Each causal LSTM contains distinct cells designed to store spatial information ($\mathcal{M}$) and temporal information ($\mathcal{C}$) from one time step to the next, as well as hidden states ($\mathcal{H}$), which are used to make predictions (Figure S2). Within the PredRNN++, which contains multiple causal LSTMs layers, each layer has its own unique temporal memory ($\mathcal{C}$) and hidden states ($\mathcal{H}$). The spatial memory ($\mathcal{M}$) is shared across all layers passed sequentially between them within a single time step. To allow the model to learn long-term dependencies, the GHU layer has its own memory cell ($\mathcal{Z}$), which bypasses the deeper causal LSTM layers and provides an alternate route for information to flow through the network. Given that the original PredRNN++ architecture was designed to handle two-dimensional inputs, we modify the implementation to handle three-dimensional inputs by replacing all two-dimensional convolutional layers with three-dimensional convolutional layers. Detailed descriptions of causal LSTMs and gradient highway unit can be found elsewhere (Wang et al., 2018).

### v2

@@ Abstract
Process-based hydrologic models are invaluable tools for understanding the terrestrial water cycle and addressing modern water resources problems. However, many hydrologic models are computationally expensive and, depending on the resolution and scale, simulations can take on the order of hours to days to complete. While techniques such as uncertainty quantification and optimization have become valuable tools for supporting management decisions, these analyses typically require hundreds of model simulations, which are too computationally expensive to perform with a process-based hydrologic model. To address this gap, we assess a hybrid modeling workflow in which a process-based model is used to generate an initial set of simulations and a machine learning (ML) surrogate model is then trained to perform the remaining simulations required for downstream analysis. As a case study, we apply this workflow to simulations of variably saturated groundwater flow at a prospective managed aquifer recharge (MAR) site. We compare the accuracy and computational efficiency of several ML architectures, including deep convolutional networks, recurrent neural networks, vision transformers, and networks with Fourier transforms. Our results demonstrate that ML surrogate models can achieve under 10% mean absolute percentage error and yield order-of-magnitude runtime savings over process-based models. Building on these findings, we examine the impacts of key modeling choices on surrogate model accuracy and efficiency. Results show that a normalized loss function improves training stability, while min-max data normalization can significantly reduce error up to a factor of 10 when compared to other treatments such as Z-score and no normalization. Downsampling input features using an autoencoder also decreases memory requirements by training with tensors 4% their original size. By reducing computational costs and memory demands, ML surrogates offer a practical pathway to scale up complex hydrologic analyses in support of sustainable groundwater management.
Note
This is the accepted version of a manuscript published in the Journal of Hydrology. The version of record is: Dai, T., Maher, K., and Perzan, Z. (2025). Machine learning surrogates for efficient hydrologic modeling: Insights from stochastic simulations of managed aquifer recharge. Journal of Hydrology, 652, 132606. doi:10.1016/j.jhydrol.2024.132606
@@ 1 Introduction
Future water security hinges on effective water resources management. Process-based hydrologic and reactive transport models, which simulate water flow and solute transport (Xu and Singh, 2004; Devia et al., 2015; Steefel et al., 2015), are frequently used to forecast future changes in the hydrologic cycle and to optimize water management strategies. When paired with uncertainty quantification and parameter estimation methods, these models can support decision-making in diverse scenarios such as flood forecasting, contaminant transport, nutrient availability, and managed aquifer recharge (MAR). These techniques often require hundreds of model simulations in which uncertain input parameters are varied randomly from one simulation to the next. While each individual simulation is deterministic, producing one fixed output for a given set of input parameters, the ensemble of simulations is stochastic because of the random variations in model input.
However, process-based simulations can be computationally expensive to perform. In particular, integrated hydrologic models — codes that simulate coupled surface and subsurface water flow — pose several numerical challenges (Maxwell et al., 2014), such as solving highly nonlinear governing equations and ensuring flux and pressure continuity at the interface between the surface and subsurface (Kollet and Maxwell, 2006). In addition, to perform simulations at the scale and resolution required for locally relevant policy decisions, model domains must resolve fine-scale heterogeneity in hydraulic conductivity, topography, soil texture, and other parameters (Wood et al., 2011). These sharp contrasts often result in steep pressure gradients that can cause iterative numerical solvers to slow down or fail to converge. As a result, performing stochastic simulations with integrated hydrologic codes requires large computational resources, even on state-of-the-art supercomputers.
@@
While these early studies highlight the potential for ML surrogate models to accelerate hydrologic simulations, key knowledge gaps remain in their application to water resources problems. First, the number of process-based simulations required to train a high-fidelity surrogate model remains unclear. Second, when generating a small number of simulations, the added cost of training an ML model may outweigh any speed gained by the trained surrogate, making it more efficient to perform simulations exclusively with a process-based model. Thus, the threshold number of simulations beyond which a hybrid workflow becomes more efficient than process-based modeling, in terms of both computational time and memory, remains poorly understood. Third, the impacts of key modeling choices, such as loss functions and data normalization techniques, on the accuracy and robustness of ML surrogates have yet to be systematically evaluated. Importantly, each of these knowledge gaps may vary across different surrogate model architectures.
In this paper, we investigate these three knowledge gaps and present practical considerations for using ML surrogates to perform stochastic hydrologic simulations. As a case study, we focus on 3D simulations of variably saturated groundwater flow during managed aquifer recharge (MAR), the intentional replenishment of depleted aquifers through recharge basins and injection wells. We train seven different surrogate model architectures to reproduce output from a parallel, integrated hydrologic code (ParFlow-CLM). We then investigate the impacts of training set size, data preprocessing techniques, and choice of loss function on surrogate model performance. We also evaluate the runtime efficiency of each surrogate architecture as compared to the process-based hydrologic code and examine techniques to increase memory efficiency, such as temporal and spatial downsampling of high-dimensional inputs using an autoencoder.
@@ 2 Material and Methods
In this section, we formalize the hybrid modeling workflow (§2.1) and introduce the case study (§2.2). Next, we describe the process-based hydrologic code (§2.3) and the surrogate model architectures (§2.4). We then describe data preprocessing (§2.5) and detail the experimental setup for training and evaluating each ML surrogate (§2.6). We conclude with a summary of experiments, enumerating the ML models trained to evaluate the effects of data normalization, loss function, and training set size on hydrologic surrogate model accuracy and the trade-offs in runtime (§2.7).
2.1 The hybrid workflow
The goal of this hybrid workflow is to harness the fast output rate of ML models to efficiently generate a large volume of simulations for downstream hydrologic analysis. The task is to execute $N$ simulations, generating $N$ simulation outputs from $N$ distinct sets of simulation inputs. As illustrated in Figure 1, we employ a traditional process-based hydrologic model (ParFlow-CLM) to generate a training set $\mathcal{D}_{\text{train}}$ comprising simulation input–output pairs. Subsequently, an ML surrogate model is trained on these examples and used to generate simulation outputs for the remaining $N-|\mathcal{D}_{\text{train}}|$ simulations, where $|\mathcal{D}_{\text{train}}|$ is the number of examples in $\mathcal{D}_{\text{train}}$. In sum, this framework produces $N$ simulations by generating $|\mathcal{D}_{\text{train}}|$ simulations with a traditional process-based hydrologic model and $N-|\mathcal{D}_{\text{train}}|$ simulations with an ML model. In a hydrologic context, these simulation outputs could aid in a multitude of downstream analyses, including uncertainty quantification, parameter estimation, and resource optimization.
@@ 2.2 The case study: Managed aquifer recharge site
We apply this hybrid workflow to a single case study in which hydrologic simulations are used to evaluate recharge efficiency at a prospective managed aquifer recharge (MAR) site. MAR is a water management strategy whereby land managers deliberately flood parcels of land or inject water into wells to replenish underlying aquifers. Recent work has shown that MAR is a valuable tool for mitigating groundwater depletion in many regions of the globe (Dillon et al., 2019; Sprenger et al., 2017), but evaluating prospective recharge sites remains challenging as subsurface heterogeneity can significantly impact recharge rates (Knight et al., 2022; Perzan et al., 2023a). Hydrologic and reactive transport simulations are effective tools for quantifying outcomes of interest at prospective MAR sites, such as changes in groundwater storage (e.g., Ganot et al., 2018) or the fate of subsurface contaminants (e.g., Perzan and Maher, 2024). However, generating these simulations comes with a notable computational expense, as the time required for a single simulation can range from hours to days depending on model scale and resolution. Thus, MAR site selection is a logical case study that could benefit from a computationally efficient hybrid modeling workflow.
@@
The domain is parameterized through a stochastic geophysical-geostatistical workflow (Perzan et al., 2023a). Because subsurface heterogeneity can exert a dominant control on recharge efficiency, we parameterize the domain using a subsurface electrical resistivity model derived from a towed transient electromagnetic (tTEM) survey of the site. The tTEM system is a hydrogeophysical tool that can map meter-scale variations in electrical resistivity, a property directly related to sediment lithology. We first generate 600 gridded realizations of the tTEM-derived resistivity model using sequential Gaussian simulation, a stochastic technique for populating a rectilinear grid with a random variable. We then compare vertical profiles of electrical resistivity to collocated profiles of sediment type collected through cone penetration tests. Using the bootstrapping procedure of Knight et al. (2018), we transform each resistivity realization to a realization of coarse fraction, a metric that describes the fraction of coarse-grained (sand and gravel) sediment within each grid cell. The remainder of sediment within the cell is assumed to be fine-grained (silt and clay). For example, a cell with a coarse fraction of 0.5 is 50% coarse-grained material (either 100% sand, 100% gravel, or some combination of the two) and 50% fine-grained material (some combination of silt and clay). To assign hydrogeologic parameters (hydraulic conductivity, porosity, etc.) to each coarse fraction realization, we assume that only two types of sediment exist within a single realization — coarse-grained and fine-grained — and that each cell is a mix of the two types. We then sample from broad distributions to assign parameter values to each end member and use a weighted averaging procedure to calculate values for cells that are a mix of the two end members. The input parameter distributions are derived from a mix of literature and site-specific data. For more information about these parameter values, we refer the reader to the more complete description provided by Perzan et al. (2023a).
@@ 2.3 The high-fidelity process-based model: ParFlow-CLM
In the first phase of our proposed framework (§2.1), a process-based hydrologic model is used to generate an initial set of simulations, which are then used to train an ML surrogate model. In this case study, we perform these process-based simulations using ParFlow-CLM. ParFlow-CLM couples the integrated hydrologic code ParFlow with the Common Land Model (CLM) to solve partial differential equations that describe water and energy fluxes over and beneath the land surface. ParFlow solves the Richards’ equation for variably saturated subsurface flow using a Newton-Krylov approach and the shallow-wave equations for overland flow using a kinematic wave approximation. CLM, meanwhile, uses a mass transfer approach to compute water and energy fluxes at the land surface (Maxwell and Miller, 2005; Kollet and Maxwell, 2008), including irrigation, evapotranspiration, and precipitation. To accurately simulate the impacts of diurnal temperature and radiation fluctuations on water fluxes across the land surface, we run ParFlow-CLM using an hourly time step in Stages 2 and 3 (§2.2). For a detailed description of the governing equations, we refer readers to previous work describing ParFlow-CLM (Jones and Woodward, 2001; Ashby and Falgout, 1996; Kollet and Maxwell, 2006; Maxwell, 2013; Maxwell and Miller, 2005; Kollet and Maxwell, 2008).
@@ 2.4 The machine learning models
In the second phase of this hybrid framework (§2.1), we train an ML surrogate on the initial set of simulations generated by the process-based model. In total, we evaluate the performance of seven different machine learning surrogate model architectures: three- and four-dimensional convolutional neural networks, three- and four-dimensional vision transformers (Dosovitskiy et al., 2021), three- and four-dimensional U-net Fourier neural operators (Wen et al., 2022; Zhou et al., 2022; Kang et al., 2023), and a three-dimensional recurrent and highway network developed for spatiotemporal forecasting, the PredRNN++ (Wang et al., 2018). All seven surrogate architectures are outlined in Table 1. We selected these architectures because some of them have been successfully applied in previous hydrologic surrogate modeling studies (Tran et al., 2021; Maxwell et al., 2021; Wen et al., 2022), while others are some of the most common architectures used for spatiotemporal forecasting in related fields like computer vision.
The ML surrogate architectures can be categorized into three types based on their method of prediction (Table 1, Figure 2). First, the Non-sequential models emulate the entire simulation in a single forward pass (i.e., the emulator produces 3D pressure fields for all output time steps in one run of the model). By contrast, the One-step models emulate one output time step (producing a single 3D pressure field) per forward pass. These models then use the output from the previous time step to perform another forward pass, repeating until all 3D pressure fields have been generated. Finally, the Recurrent models are similar to the one-step models in that they emulate one output time step per forward pass. However, they use memory cells to retain hidden states between forward passes, so that information can be stored between output time steps. As discussed above, the output frequency varies between model stages (see 2.2), which results in a different shape input tensor for each model stage. Because most ML architectures require a consistent shape input tensor across all training examples, we train a separate surrogate for each model stage. The ML models are summarized below, though complete details are provided in the Supplementary material (Section S1.1).
@@
We experiment with three non-sequential models. Since non-sequential models produce 3D pressure fields for all output time steps in a single forward past, they are four-dimensional; i.e., they receive four-dimensional inputs — x, y, z, and time — and generate four-dimensional outputs (3D transient pressure fields). The non-sequential models are:
@@
    ViT4d. The ViT4d is a vision transformer (Dosovitskiy et al., 2021) that uses a four-dimensional convolutional layer to transform the input tensor into a sequence of four-dimensional patches. This sequence of patches is then fed into a self-attention module (Vaswani et al., 2017) that calculates the importance of each patch relative to every other patch, allowing the model to learn both spatial and temporal dependencies regardless of their distance from each other.
@@
One-step models, unlike non-sequential models, receive three-dimensional inputs (x, y, and z) and predict one time step at a time. The output of the previous time step is reused as input for the next time step and this rolling forward process continues until the final time step. Note that these models receive no explicit temporal information during training, such as day of year or time step size. These one-step models include:
@@
    PredRNN++. This recurrent model is an improved predictive recurrent neural network (PredRNN++), which was originally developed by Wang et al. (2018). The PredRNN++ architecture (Figure S2) pairs causal long short-term memory networks (causal LSTMs) with a gradient highway units (GHUs). Causal LSTMs contain a spatiotemporal memory mechanism designed to improve short-term spatial correlations, while GHUs connect future outputs with distant inputs for improved long-term modeling capabilities. Each causal LSTM contains distinct cells designed to store spatial information ($\mathcal{M}$) and temporal information ($\mathcal{C}$) from one time step to the next, as well as hidden states ($\mathcal{H}$), which are used to make predictions (Figure S2). Within the PredRNN++, which contains multiple causal LSTM layers, each layer has its own unique temporal memory ($\mathcal{C}$) and hidden states ($\mathcal{H}$). The spatial memory ($\mathcal{M}$) is shared across all layers passed sequentially between them within a single time step. To allow the model to learn long-term dependencies, the GHU layer, equipped with its own memory cell ($\mathcal{Z}$), is called once per time step, between the first and second causal LSTM layers, and functions similarly to a residual connection. The GHU allows information and gradients to bypass deep-in-time model layers, reducing the risk of vanishing gradients that result from long-term predictions. Given that the original PredRNN++ architecture was designed to handle two-dimensional inputs, we modify the implementation to handle three-dimensional inputs by replacing all two-dimensional convolutional layers with three-dimensional convolutional layers. Detailed descriptions of causal LSTMs and gradient highway unit can be found elsewhere (Wang et al., 2018).